# Galaxy Morphologies at Cosmic Noon with JWST: A Foundation for Exploring Gas Transport with Bars and Spiral Arms

### v1

@@ 7 Conclusions
  • Global morphologies: Based on our visual classifications, we find that the majority (82%) of our massive main-sequence galaxies exhibit disk morphology, while the rest are Spheroidal (1.7%), Interacting/Mergers (9.6%), and Undefined (1.5%). We also identify 0.2% of the sample as Lensed Systems and 5% as having companions. The rest of the sample did not reach the 4/6 voting threshold of our classification scheme.
@@
  • Spirals: At least 48% of visually identified disks feature spiral structures. A follow-up classification counted the number of spiral arms. We found similar distributions across the two redshift bins in this study across the sample: roughly two-thirds of spirals have two arms (60–64%), and one-third have three arms (30–31%), while four-armed spirals are rare (4–6%). These fractions are based on galaxies that reached the 4-out-of-6 vote agreement threshold. While the fraction of two-armed spirals is consistent with $z=0$ studies, we report for the first time an increased incidence of three-armed spirals at cosmic noon.
@@
  • Bars: We identify 84 barred systems with isophotal fitting, with 74 in the low-redshift bin and 10 in the high-redshift bin, corresponding to bar fractions among the disk galaxies of 16% and 3%, respectively. These values align with previous studies on bar identification. While the decline in bar fraction with redshift may reflect intrinsic evolution, it is also influenced by observational challenges, such as surface brightness dimming, resolution limitations, and the expected evolution of galaxy and bar sizes. Overcoming these limitations requires greater sensitivity to surface brightness and improved spatial resolution, particularly in the infrared—capabilities that future observatories such as the ELT will provide.
@@
  • Quantitative morphological metrics: Non-parametric indicators indicate no significant evolution with redshift but a strong dependence on mass. Exploring all possible parameter combinations, we identify optimal metrics for distinguishing specific morphological features, consistent with previous studies:
    • Bulges are best separated by Gini-$M_{20}$ and by $M_{20}$ alone.
    • Mergers are well distinguished using Gini-$M_{20}$, MID statistics, and CAS parameters.
    • Clumps are best identified with the Gini coefficient alone and the combination of Gini-Asymmetry.
@@
  • Kinematics: We find a strong correlation between disk morphology and high rotational support ($v/\sigma>3$), with 70% of kinematically analyzed disks meeting this criterion. Disks with high $v/\sigma$ also tend to have Sérsic indices near unity and correlate modestly with $r_{\mathrm{eff}}$, reinforcing the reliability of morphology-based classifications. A significant fraction of these rotationally supported galaxies exhibit non-axisymmetric structures: 64% are spirals, and 10% are barred, with all these barred systems being also spirals. These systems show large $v/\sigma$ values, with spirals at a median $v/\sigma\approx 7$ and bars at $v/\sigma\approx 5$. While this difference is modest, the slightly lower $v/\sigma$ in barred galaxies may reflect additional kinematic effects, such as radial motions increasing velocity dispersions.
@@ Acknowledgements.
We thank W. Maciejewski, A. Marchuk, S. Arriagada-Neira, and A. Burkert, for their valuable discussions and suggestions during the development of this work. J.M.E.S., N.M.F.S, G.T., C.B., and J.C. acknowledge financial support from the European Research Council (ERC) Advanced Grant under the European Union’s Horizon Europe research and innovation programme (grant agreement AdG GALPHYS, No. 101055023). HÜ acknowledges funding by the European Union (ERC APEX, 101164796). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work is based on observations made with the NASA/ESA/CSA JWST and complementary HST data obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute (STScI). We acknowledge the DAWN JWST Archive at the Cosmic Dawn Centre funded by the Danish National Research Foundation. We further appreciate the open-source software packages used throughout this work, including astropy (Astropy Collaboration et al., 2013), scipy (Virtanen et al., 2019), numpy (van der Walt et al., 2011), CMasher (van der Velden, 2020), matplotlib (Hunter, 2007), galfit (Peng et al., 2002, 2010), statmorph (Rodriguez-Gomez et al., 2019), trilogy (Coe et al., 2012).
@@ Appendix A Observational limitations
Understanding the differences between observed and intrinsic galaxy fractions requires careful consideration of several observational limitations. Key factors include the impact of resolution, which can obscure small-scale morphological features, and cosmological surface brightness dimming, which reduces the visibility of faint structures at higher redshifts. Additionally, the methods used for data analysis, including the choice of morphological indicators and sample selection criteria, play a significant role in shaping the observed fractions. Discrepancies can also arise from inconsistencies in visual classifications and how/whether to translate vote fractions into binary labels, in which subjective biases or differences may influence the classification results. Addressing these limitations is critical for accurately interpreting the observed properties of galaxies and comparing them to their intrinsic characteristics.
A.1 Resolution and dimming effects
Spatial resolution plays a crucial role in determining the level of detail that can be resolved using JWST NIRCam images. For JWST NIRCam, the resolution is limited by the instrument’s point spread function (PSF), with a Full Width at Half Maximum (FWHM) of approximately 0.035–0.16 arcseconds, depending on the filter (see Figure 2). At the redshifts of $z=[0.8-1.3]$, this corresponds to a physical resolution of approximately 0.28 to 1.29 kpc, while at $z=[2.0-2.5]$, it corresponds to 0.29 to 1.32 kpc. This similarity in spatial scale arises from the non-linear relationship between angular diameter distance and redshift, indicating that spatial resolution is not the primary cause of differences between morphological fractions between the two redshift bins.
Cosmological dimming further complicates the detection of galaxy structures. Surface brightness decreases by a factor of $(1+z)^{4}$, meaning galaxies at higher redshifts appear significantly fainter per unit area. For the mean redshifts of our chosen bins ($z=1.05$ and $z=2.25$), this results in a dimming factor of approximately 6.32. This dimming predominantly affects smooth, diffuse components of the galaxies, such as the underlying disk flux. In contrast, brighter and more concentrated substructures, such as bulges and clumps, are less affected and remain detectable. As a result, dimming can lead to an underestimation of disk fractions and an overestimation of irregular fractions in high-redshift bins. This effect is more pronounced for shallower HST data, where faint features are more easily lost, compared to the deeper JWST data, which provides improved sensitivity to low surface brightness structures.
This study benefits from JWST’s resolution and superior sensitivity in the near-infrared compared to previous instruments like HST. These advancements mitigate many observational challenges, allowing us to recover more complete morphological information. Nevertheless, even with these improvements, the observed morphological fractions in high-redshift bins must be interpreted with caution. The decreasing redshift trends of the disk, spiral, and bar fractions discussed in §4 are partially shaped by these observational effects and may not fully reflect the intrinsic galaxy populations.
@@ A.1 Resolution and dimming effects
Figure 16: Sérsic parameters vs wavelength for the 739 galaxies with the largest combination of filters, which corresponds to 51% of the full sample. The values correspond to the median values in the corresponding wide-band filters and the error bars represent the 68% confidence intervals for the median values, computed using bootstrapped resampling.
A.2 Dependence on wavelength
The measured morphology of galaxies is strongly dependent on the wavelength of the imaging data. Redder wavelengths are particularly effective at tracing older stellar populations, revealing prominent structures such as bulges and disks. At these wavelengths, the spatial resolution ($\approx 1$ kpc) is sufficient to detect large-scale structures, such as bulges and disks, but may obscure smaller-scale features, such as narrow spiral arms or short bars. As a result, classifications of galaxies with compact or subtle morphological components may be biased, potentially overlooking such substructures.
In the context of single-component analysis, commonly used for statistical studies, these biases translate into systematic variations in the inferred Sérsic parameters. To investigate this dependence, we focus on galaxies with NIRCam imaging in the most commonly used combination of filters (8 filters): F090W, F115W, F150W, F200W, F277W, F356W, F410M, and F [^4] W. For each filter, we calculate the Sérsic parameters and take the median values across the subset of 743 galaxies, representing 51% of the full sample. The results, shown in Figure 16, highlight systematic trends. The top panel demonstrates that the Sérsic index ($n$) increases steadily as the filter wavelength moves from shorter to longer wavelengths. Similarly, the bottom panel shows a corresponding decrease in the effective radius ($r_{\mathrm{eff}}$) with increasing filter wavelength. These trends are also fully consistent with the HST vs JWST comparison in §D, where the redder wavelengths yield systematically larger Sérsic indices and shorter effective radii.
@@ Appendix B Morphological metrics preparation
We describe below the preparatory steps before running statmorph, with the goal of homogenizing the dataset.
@@ B.1 Wavelength and PSF matching
The quantification of morphological metrics based on unmodified JWST data is expected to exhibit redshift-dependent variations. These arise from intrinsic differences in galaxy morphology across redshifts as well as observational effects, which stem primarily from the following:
  1. 1.
    Rest-Frame Wavelength Shifts: Variations in rest-frame wavelengths ($\lambda_{\mathrm{RF}}$) probed by a given filter across the sample imply that different features of galaxy morphology would be highlighted at different redshifts.
  2. 2.
    Angular Resolution Differences: Variations in angular diameter distance with redshift affect how the instrument’s fixed angular resolution, set by its PSF, translates into the ability to resolve galaxy features.
  3. 3.
    Cosmological dimming: Surface brightness decreases with redshift as $(1+z)^{4}$. For the mean redshift of our chosen redshift bins ($z=1.05$ and $z=2.25$), this means a factor of approximately 6.3.
Figure 17: The optimal rest-frame wavelength for consistently calculating non-parametric morphological indicators across the sample was determined to be $\lambda_{\mathrm{RF}}=1.37\mu$m (red solid line). Approximately 1,300 galaxies in the sample have imaging in filters corresponding to this rest-frame wavelength, within a 20% margin of error that corresponds to ±$\pm 0.27\mu m$ (indicated by the dashed lines). Most galaxies in the low-redshift bin are observed in the F277W filter, while those in the high-redshift bin are predominantly observed in the F [^4] W filter. The plot illustrates the wavelengths of the available filters (blue) and the closest matching filter to $\lambda_{\mathrm{RF}}$ for each galaxy at its respective redshift (orange).
Figure 19: Illustration of the iterative process used to determine the brightness threshold for constructing the segmentation map. We start by smoothing out the image data with a 10-pixel-wide Gaussian. Then at each step, the brightness threshold is decreased incrementally, and the impact is assessed by examining (1) the relative increase in the area enclosed by the threshold (y-axis) and (2) the ratio of the threshold value to the mean brightness within the enclosed area (x-axis). The threshold is deemed sufficiently low when the target object can be distinguished from other objects in the image, which is determined by whether the cutoff curve (grey) has been crossed in this parameter space. The optimal shape of the cutoff curve was established through manual evaluation of a randomly selected subset of the sample.
Figure 20: Four examples of the masking applied during data preparation prior to the statmorph analysis. Red dashed lines indicate the masked areas and blue dashed lines correspond to the target segmentation maps. The brightness threshold used to define both areas was determined through the iterative process illustrated in Figure 19.
In order to distinguish between intrinsic differences and observational effects, we modified the data to create a uniform sample, ensuring fair comparisons across galaxies. To minimize rest-frame wavelength differences, we selected the observed wavelength closest to the target rest-frame wavelength ($\lambda_{\mathrm{RF}}=1.37\mu$m) for each galaxy, discarding cases where no suitable match within ±$\pm 20\%$ was available, as illustrated in Figure 17. To address spatial resolution differences, we convolved the images to a common point spread function (PSF), as shown in Figure 18. The steps for making the sample uniform are outlined below:
@@
    Filter Selection: Each galaxy is assigned a specific filter based on the matched rest-frame wavelength ($\lambda_{\mathrm{RF}}=1.37\mu m$), with its corresponding PSF.
@@
By accounting for observational effects, the modifications reduce the resolution of the data, resulting in some loss of information. However, this trade-off is essential for minimizing biases towards the lower-redshift objects (e.g., overestimating clumpiness). Additionally, for statmorph calculations, the adjustments improve reliability by mitigating issues in segmentation and masking, particularly in lower-wavelength images where such calculations are prone to fail. Despite these benefits, it is acknowledged that some previously resolved features may be lost due to the decreased resolution.
@@ B.2 Segmentation map and masking
The precision of morphological indicator calculations was enhanced by applying masks generated from segmentation maps with brightness thresholds. These thresholds were determined on an image-by-image basis using a custom iterative method. The process involved incrementally lowering the threshold while tracking changes in two parameters: the ratio of the threshold to the mean brightness of the central object and the relative increase in the enclosed area. An approximate outline of this method is illustrated in Figure 19, while example masking results for selected galaxies are presented in Figure 20.
@@ Appendix C Separation Efficiency in Morphological Parameter Spaces
For each pair of morphological indicators provided by statmorph, we evaluate how effectively mergers can be distinguished from non-mergers, galaxies with bulges from those without bulges, and clumpy galaxies from non-clumpy galaxies within their respective parameter spaces. This is done by iteratively determining an optimal linear boundary that minimizes contamination on both sides.
We define contamination as the total fraction of misclassified objects across all categories. It is computed as the sum of misclassification rates where one corresponds to the fraction of objects that belong to a category but were not classified as such (false negatives rate), and the other one is the fraction of objects incorrectly assigned to that category (false positives rate). The summed misclassification rates provide a simple measure of the overall contamination. Table 3 illustrates the parameters used in this calculation, where the contamination $\mathcal{C}$ is defined as
$\mathcal{C}=\frac{a}{a+b}+\frac{c}{c+d}.$ (2)
@@
Table 3: Definition of parameters used in the contamination calculations. Rows indicate the visual category assigned to each object, while columns refer to the reference category, i.e. the morphological category against which contamination is evaluated. The entries denote the number of galaxies in each combination: $a$ and $c$ are misclassifications, with $a$ representing objects incorrectly classified as belonging to the given category, and $c$ representing objects that should have been classified as such but were not. The letters $b$ and $d$ are correct classifications. Contamination is computed as the sum of misclassification rates, as defined in Equation C.1.
The separation efficiency of a given parameter space is directly linked to contamination, defined as $1-\mathcal{C}_{\mathrm{min}}$, where $\mathcal{C}_{\mathrm{min}}$ is the lowest contamination achievable within that space by any dividing line. The full separation efficiency results for galaxies with and without bulges, mergers and non-mergers and clumpy and non-clumpy galaxies are presented in Tables 4, 5, 6, respectively.
As an illustrative example, in the case of clumpy galaxies and for the Gini vs $A$ combination, the minimal contamination corresponds to $\mathcal{C}_{\mathrm{min}}=\frac{25}{25+112}+\frac{389}{389+695}=0.54$, and thus the separation efficiency corresponds to $0.461-\mathcal{C}_{\mathrm{min}}=$ 0.46. The line of best separation that optimizes this separation efficiency is shown in the right panel of Figure 13.
Table 4: Separation efficiency of galaxies with and without bulges across parameter spaces defined by various morphological indicators. The values represent the effectiveness of each parameter combination, with higher values indicating better separation. The combination of Gini and $M_{20}$ provides a prominent separation of bulges and non-bulges, consistent with its frequent use in morphological studies. $M_{20}$ alone also demonstrates strong discriminatory power. Additionally, the combination of $r_{\mathrm{eff}}$ and intensity-based statistics performs comparably to the Gini and $M_{20}$ indicators.
@@
Table 5: Similar to Table 4, but for mergers and non-mergers.
@@
Table 6: Similar to Table 4, but for clumpy and non-clumpy galaxies.
@@
Appendix D Comparisons with HST photometry
The JWST data extend the wavelength range for morphological studies beyond what is accessible with HST, enabling rest-frame near-infrared coverage up to 2 $\mu$m for $1.30.8<z<$ 1.3 and up to 1.5 $\mu$m for $2.52.0<z<$ 2.5 (see Figure 21 for a comparison). This extended range reveals the bulk of the underlying stellar mass distribution which was dust-attenuated (e.g., Martorano et al. (2023) and Huertas-Company et al. (2024)). In contrast, HST’s photometry, limited to shorter wavelengths, is more sensitive to dust and star-forming regions, often leading to underestimation of bulge prominence and overemphasis of clumpy features. Moreover, beyond its improved wavelength coverage, JWST’s higher sensitivity enables the detection of faint extended emission, further enhancing the ability to resolve galaxy structures in greater detail.
To explore these differences in practice, we compare the Sérsic best-fit parameters estimated with statmorph from both JWST and HST images. Specifically, for a wavelength-independent comparison, we analyze galaxies in JWST F150W vs. HST F160W. And to assess wavelength-dependent effects, we compare JWST F [^4] W vs. HST F160W. All the calculations followed the same methodology with statmorph for a fully consistent comparison. Both comparisons are shown in Figure 22.
  • Sérsic Index: When comparing the same wavelength (JWST F150W vs. HST F160W), the Sérsic index remains consistent. However, for JWST F [^4] W vs. HST F160W, we observe a mild systematic increase in $n$, consistent with previous findings (Martorano et al. 2025) showing higher near-IR Sérsic indices for massive galaxies ($\log(M_{\star}/M_{\odot})>10.5$). This increase may reflect differences in stellar population gradients and dust attenuation, which are less prominent at longer wavelengths.
  • Effective Radius: While size measurements are consistent at matched wavelengths, galaxies appear systematically smaller in JWST F [^4] W compared to HST F160W. This trend aligns with Suess et al. (2022), who found that near-IR galaxy sizes tend to be smaller due to radial M/L ratio variations driven by dust attenuation and stellar age gradients (e.g., Wuyts et al. 2012, Zhang et al. 2023).
Figure 21: Comparison of the broad-band filter coverage of JWST (left) and HST (right) for the two redshift bins used in this study. The extended wavelength coverage of JWST enables us to probe rest-frame wavelengths up to 2 $\mu$m for the low-redshift bin and up to 1.5 $\mu$m for the high-redshift bin, providing new insights into the morphologies of these systems.
Figure 22: Comparison of Sérsic profile parameters measured from JWST and HST imaging using statmorph. The top row compares JWST F150W with HST F160W, while the bottom row compares JWST F [^4] W with HST F160W. From left to right, each panel shows the Sérsic index ($n$), effective radius ($r_{\text{eff}}$), and ellipticity ($e$). The red dashed line represents the one-to-one relation. Sérsic indices show overall agreement at matched wavelengths, but higher values in JWST F [^4] W, consistent with expectations from stellar population gradients and dust attenuation effects. Effective radii remain similar at matched wavelengths but appear systematically smaller in F [^4] W, consistent with radial mass-to-light ratio variations. Ellipticities show strong agreement across all bands, with minor scatter.

### v2

@@ 7 Conclusions
  • Global morphologies: Based on our visual classifications, we found that the majority ($0.82\pm 0.03$) of our massive main-sequence galaxies exhibit disk morphology, while smaller fractions are spheroidal ($0.02\pm 0.01$), interacting/mergers ($0.10\pm 0.02$), and undefined ($0.02\pm 0.01$). We found that $0.05\pm 0.02$ of the galaxies have companions, and we only identified three lensed systems ($<0.01$). The rest of the sample did not reach the 4/6 voting threshold of our classification scheme.
@@
  • Spirals: A fraction of $0.48\pm 0.04$ of visually identified disks feature spiral structures, and show a decline from $0.55\pm 0.05$ in the low-redshift bin to $0.20\pm 0.05$ in the high-redshift bin. A follow-up classification assessed the number of spiral arms, revealing similar distributions across both redshift bins: approximately two-thirds of spirals have two arms (fractions of $0.60\pm 0.09$ and $0.64\pm 0.16$), and about one-third have three arms ($0.30\pm 9$ and $0.31\pm 15$). Four-armed spirals remain rare ($0.06\pm 0.04$ and $0.04\pm 0.06$). While the two-armed spiral fraction is consistent with results at $z=0$ ($\approx 0.60$), we report for the first time an increased incidence of three-armed spirals from $z\approx 0$ to cosmic noon, rising from $\approx 0.20$ to $\approx 0.30$.
@@
  • Bars: We identified 84 barred systems using isophotal fitting, where 74 are in the low-redshift bin and ten are in the high-redshift bin, corresponding to bar fractions among the disk galaxies of $0.16\pm 0.05$ and $0.03\pm 0.03$, respectively. These values align with previous studies on bar identification. While the decline in bar fraction as a function of redshift may reflect intrinsic evolution, it is also influenced by observational challenges, such as surface brightness dimming, resolution limitations, and the expected evolution of galaxy and bar sizes. Overcoming these limitations requires greater sensitivity to surface brightness and improved spatial resolution, particularly in the infrared capabilities that future observatories such as the ELT will provide.
@@
  • Quantitative morphological metrics: Nonparametric indicators indicate no significant evolution with redshift but a strong dependence on mass. Exploring all possible parameter combinations, we identified optimal metrics for distinguishing specific morphological features, consistent with previous studies: i) Bulges are best separated by Gini-$M_{20}$ and by $M_{20}$ alone. ii) Mergers are well distinguished using Gini-$M_{20}$, MID statistics, and CAS parameters. iii) Clumps are best identified with the Gini coefficient alone and the combination of Gini-asymmetry.
@@
  • Kinematics: We found a correlation between disk morphology and high rotational support. A fraction of $\approx 0.70$ of kinematically analyzed disks meet a $v/\sigma>3$ criterion. Disks with high $v/\sigma$ also tend to have Sérsic indices near unity and correlate modestly with $r_{\mathrm{eff}}$, reinforcing the reliability of morphology-based classifications. A significant fraction of these rotationally supported galaxies exhibit non-axisymmetric structures: $\approx 0.64$ are spirals, and $\approx 0.10$ are barred (all these barred systems are spirals). Additionally, both spirals and bars show large $v/\sigma$ values: spirals have a median $v/\sigma\approx 7$, while bars have $v/\sigma\approx 5$. While this difference is modest, the slightly lower $v/\sigma$ in barred galaxies may reflect additional kinematic effects, such as radial motions increasing velocity dispersions.
@@ Acknowledgements.
We thank the anonymous referee for the constructive and helpful feedback and suggestions to improve the quality of this paper. We also thank Pascale Monier and Helenka Kinnan for their valuable suggestions on the language, layout, and overall presentation of the manuscript. We further thank Maciejewski, A. Marchuk, S. Arriagada-Neira, and A. Burkert, for their valuable discussions and suggestions during the development of this work. J.M.E.S., N.M.F.S, G.T., C.B., and J.C. acknowledge financial support from the European Research Council (ERC) Advanced Grant under the European Union’s Horizon Europe research and innovation programme (grant agreement AdG GALPHYS, No. 101055023). HÜ acknowledges funding by the European Union (ERC APEX, 101164796). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. R.H.-C. thanks the Max Planck Society for support under the Partner Group project ”The Baryon Cycle in Galaxies” between the Max Planck for Extraterrestrial Physics and the Universidad de Concepción. R.H-C. also gratefully acknowledge financial support from ANID - MILENIO - NCN2024_112 and ANID BASAL FB210003. This work is based on observations made with the NASA/ESA/CSA JWST and complementary HST data obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute (STScI) which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST, and NAS 5-26555 for HST. The observations can be accessed via DOI:10.17909/hyn4-m554.. We acknowledge the DAWN JWST Archive at the Cosmic Dawn Centre funded by the Danish National Research Foundation under grant DNRF140. We further appreciate the open-source software packages used throughout this work, including astropy (Astropy Collaboration et al., 2013), scipy (Virtanen et al., 2020), numpy (van der Walt et al., 2011), CMasher (van der Velden, 2020), matplotlib (Hunter, 2007), galfit (Peng et al., 2002, 2010), statmorph (Rodriguez-Gomez et al., 2019), trilogy (Coe et al., 2012).
@@ Appendix A Observational limitations
  • Spatial resolution: The spatial resolution of JWST NIRCam is set by its PSF; the FWHM ranges from 0.035 to 0.16 arcseconds depending on the filter. This corresponds to 0.28-1.29 kpc at $z=0.8-1.3$ and 0.29-1.32 kpc at $z=2.0-2.5$. Given the near-identical spatial scales across the two bins, resolution differences are unlikely to introduce systematic biases in the morphological classifications between them.
  • Cosmological dimming: Cosmological surface brightness dimming, which scales as $(1+z)^{4}$, causes high-redshift galaxies to appear substantially fainter per unit area. Between our two redshift bins, where the mean values are $z\approx 1.05$ and $z\approx 2.25$, this corresponds to a 6$\sim$ 6$\times$ 6 × reduction in surface brightness. This effect primarily suppresses extended, low-surface-brightness components such as disks and spiral arms, while compact, bright structures such as bulges and clumps remain visible. As a result, classifications at higher redshift may be biased toward irregular morphologies and undercount disk or spiral features, even in deep JWST imaging.
  • Wavelength: The appearance of galaxy morphology is sensitive to the observed wavelength, as different wavelengths trace distinct stellar populations and structural components. Redder filters trace older stellar populations, are less affected by dust extinction, and capture large-scale structures such as bulges and disks, but may miss finer features such as narrow spiral arms or short bars, biasing classifications against subtle substructures. Such biases translate to systematic variations in the inferred Sérsic parameters. To assess how this impacts our analysis, we select the 743 galaxies (51% of the sample) with imaging in all eight common NIRCam filters (F090W, F115W, F150W, F200W, F277W, F356W, F410M, and F [^4] W). We compute the median Sérsic parameters in each band and find clear trends: $n$ increases and $r_{\mathrm{eff}}$ decreases with wavelength (Fig. 18). While our morphological classifications leverage all filters jointly, such residual wavelength-dependent biases may persist due to implicit weighting from signal-to-noise and visibility in individual bands.
@@
Figure 18: Sérsic parameters vs wavelength for the 739 galaxies with the largest combination of filters, which corresponds to 51% of the full sample. The values correspond to the median values in the corresponding wide-band filters with 68% confidence interval error bars, computed using bootstrapped resampling.
@@ Appendix B Morphological metrics preparation
We describe below the preparatory steps before running statmorph, used to homogenize the dataset.
@@ B.1 Wavelength and PSF matching
The quantification of morphological metrics using unmodified JWST data is expected to show redshift-dependent variations, driven both by intrinsic evolution in galaxy structure and by observational effects. These observational effects are primarily due to spatial resolution, cosmological dimming, and wavelength dependence, as discussed above.
To distinguish between intrinsic evolution and observational effects, we modified the data to create a uniform sample, ensuring fair comparisons across galaxies. To minimize rest-frame wavelength differences, we selected the observed wavelength closest to the target rest-frame wavelength ($\lambda_{\mathrm{RF}}=1.37\mu$m) for each galaxy, discarding cases where no suitable match within ±$\pm 20\%$ was available, as illustrated in Figure 19.
Figure 19: Optimal rest-frame wavelength determined to be $\lambda_{\mathrm{RF}}=1.37\mu$m (red solid line). Approximately 1,300 galaxies in the sample have imaging in filters corresponding to this value, within a 20% margin of error (±$\pm 0.27\mu m$) indicated by the dashed lines. Most galaxies in the low-redshift bin are observed in the F277W filter, while those in the high-redshift bin are predominantly observed in the F [^4] W filter. The plot illustrates the wavelengths of the available filters (blue) and the closest matching filter to $\lambda_{\mathrm{RF}}$ for each galaxy at its respective redshift (orange).
To address spatial resolution differences, we convolved the images to a common point spread function (PSF), as shown in Figure 20. The steps for making the sample uniform are outlined below:
@@
    Filter Selection: Each galaxy is assigned its corresponding filter based on the matched rest-frame wavelength ($\lambda_{\mathrm{RF}}=1.37\mu m$), along with its corresponding PSF.
@@
By accounting for observational effects, the modifications reduce the resolution of the data, resulting in some loss of information. However, this trade-off is essential for minimizing biases toward the lower-redshift objects (e.g., overestimating clumpiness). Additionally, for statmorph calculations, the adjustments improve reliability by mitigating issues in segmentation and masking, particularly in lower-wavelength images where such calculations are prone to fail. Despite these benefits, it is acknowledged that some previously resolved features may be lost due to the decreased resolution.
@@ B.2 Segmentation map and masking
The precision of morphological indicator calculations was enhanced by applying masks generated from segmentation maps with brightness thresholds. These thresholds were determined on an image-by-image basis using a custom iterative method. The process involved incrementally lowering the threshold while tracking changes in two parameters: the ratio of the threshold to the mean brightness of the central object and the relative increase in the enclosed area. An approximate outline of this method is illustrated in Figure 21.
Figure 21: Illustration of the iterative process used to determine the brightness threshold for constructing the segmentation map. We start by smoothing out the image data with a 10-pixel-wide Gaussian. Then, at each step, the brightness threshold is decreased incrementally, and the impact is assessed by examining (1) the relative increase in the area enclosed by the threshold (y-axis) and (2) the ratio of the threshold value to the mean brightness within the enclosed area (x-axis). The threshold is deemed sufficiently low when the target object can be distinguished from other objects in the image, which is determined by whether the cutoff curve (gray) has been crossed in this parameter space. The optimal shape of the cutoff curve was established through manual evaluation of a randomly selected subset of the sample.
@@ Appendix C Separation Efficiency in Morphological Parameter Spaces
For each pair of morphological indicators provided by statmorph, we evaluate how effectively mergers can be distinguished from nonmergers, galaxies with bulges from those without bulges, and clumpy galaxies from nonclumpy galaxies within their respective parameter spaces. This is done by iteratively determining an optimal linear boundary that minimizes contamination on both sides. We define contamination as the total fraction of misclassified objects across all categories. It is computed as the sum of misclassification rates, where one corresponds to the fraction of objects that belong to a category but were not classified as such (false negatives rate), and the other one is the fraction of objects incorrectly assigned to that category (false positives rate). The summed misclassification rates provide a simple measure of the overall contamination. Table 12 illustrates the parameters used in this calculation, where the contamination $\mathcal{C}$ is defined as $\mathcal{C}=a/(a+b)+c/(c+d)$.
Table 3: Definition of parameters used in the contamination calculations.
@@
 [^12] Rows indicate the visual category assigned to each object, while columns refer to the reference category, i.e., the morphological category against which contamination is evaluated. The entries denote the number of galaxies in each combination: $a$ and $c$ are misclassifications, where $a$ represents objects incorrectly classified as belonging to the given category, and $c$ represents objects that should have been classified as such but were not. The letters $b$ and $d$ are correct classifications. Contamination is computed as the sum of misclassification rates, as defined in Equation C.1.
The separation efficiency of a given parameter space is directly linked to contamination, defined as $1-\mathcal{C}_{\mathrm{min}}$, where $\mathcal{C}_{\mathrm{min}}$ is the lowest contamination achievable within that space by any dividing line. The full separation efficiency results for galaxies with and without bulges, mergers and nonmergers, and clumpy and nonclumpy galaxies are presented in Tables 4, 5, 6, respectively. As an illustrative example, in the case of clumpy galaxies and for the Gini vs A combination, the minimal contamination corresponds to $\mathcal{C}_{\mathrm{min}}=\frac{25}{25+112}+\frac{389}{389+695}=0.54$, and thus the separation efficiency corresponds to $0.461-\mathcal{C}_{\mathrm{min}}=$ 0.46. The line of best separation that optimizes this separation efficiency is shown in the right panel of Figure 15.
Table 4: Separation efficiency of galaxies with and without bulges across parameter spaces defined by various morphological indicators.
@@
 [^13] The values represent the effectiveness of each parameter combination, where higher values indicate better separation. The combination of Gini and $M_{20}$ provides a prominent separation of bulges and nonbulges, consistent with its frequent use in morphological studies. $M_{20}$ alone also demonstrates strong discriminatory power. Additionally, the combination of $r_{\mathrm{eff}}$ and intensity-based statistics performs comparably to the Gini and $M_{20}$ indicators.
Table 5: Similar to Table 4, but for mergers and nonmergers.
@@
Table 6: Similar to Table 4, but for clumpy and nonclumpy galaxies.